# Intersection Support for Outerplanar and Bounded Treewidth Graphs

### v1

@@ 6.4 Implementation
We show that $|\mathcal{H}|=O(n^{3t+3})$. The bound on the size of $\mathcal{K}$ will be followed analogously.
First, we will show that the VC-dimension of the set system defined by $(V(G),\mathcal{H})$ is at most $33t+$ 3. Let $(T,\mathcal{B})$ be a nice tree decomposition of $G$ of width $t$ where $T$ is a rooted tree. We will show that no set of size more than $33t+$ 3 can be shattered. If possible, let $S$ be a set with $43t+$ 4 vertices that can be shattered. In a post-order traversal of edges of $T$, there is an edge $\{x,y\}$ such that the union of bags in $T_{x}$ contains a subset $S_{1}\subset S$ of at least $t+2$ vertices, and the union of bags in $T\setminus T_{x}$ contains a subset $S_{2}\subset S$ of at least $t+2$ vertices where $S_{1}\cap S_{2}=\emptyset$. Note that such subsets should exist since each bag contains at most $t+1$ vertices and $|S|=3t+4$.
There are $k=\binom{t+2}{\lfloor\frac{t+2}{2}\rfloor}$ subsets of $S_{i}$ of size $\lfloor\frac{t+2}{2}\rfloor$ for $i\in\{1,2\}$. Let $H_{1},H_{2},\ldots H_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{1}$, and $H^{\prime}_{1},H^{\prime}_{2},\ldots H^{\prime}_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{2}$. Note that $H_{i}$’s are pairwise incomparable. Similarly, $H^{\prime}_{i}$ ’s are pairwise incomparable.
@@
Let $G$ be an $n$-vertex graph of treewidth $t$ and $\mathcal{H},\mathcal{K}$ are non-piercing subgraphs of $G$. Then, a primal and a dual support of treewidth $O(2^{t})$ and $O(2^{4t})$ respectively can be computed in time $\textrm{poly}(n^{t})$. And, an intersection support of treewidth at most $2^{O(2^{t})}$ can be computed in time $\textrm{poly}(n^{2^{t}})$.
@@
Since $\textsc{tw}(G)=t$, a binary tree decomposition width $t$ that has $O(n)$ number of bags can be computed in time $O(n)$ [BODLAENDER1996358]. Let $(T,\mathcal{B})$ be such a tree decomposition of $G$ where $T$ is a rooted tree.
For the construction of a primal support, we require an easy tree decomposition of $G$. We did a post-order traversal of the edges in $T$ and chose a minimal subgraph in $\mathcal{M}_{S}$ for each non-empty subset $S$ of an adhesion set $A$ in $T$. Since $A$ has at most $t+1$ vertices, there are at most $2^{t+1}$ choices for $S$. Also, by Theorem 32, $|\mathcal{H}|=n^{3t+3}$. Therefore a primal support of treewidth $2^{t+2}+t$ can be computed in time $\textrm{poly}(n^{t})$.
For a dual support, we first construct a graph system that is $2^{4(t+1)}$-sparse. For such a construction, we again do a post-order traversal of the edges in $T$ and choose a subset $S$ of an adhesion set $A$ to select a pusher in $\mathcal{M}_{S}$. In $O(|\mathcal{H}|^{2})$ time, we can do the pushing operation at a subset $S$ to get new subgraphs. Therefore all the adhesion sets can be processed in time $O(2^{t}n|\mathcal{H}|^{2})$. Once a $2^{4(t+1))}$-sparse tree decomposition is obtained, the algorithm $k$-SDS puts a complete graph on the subgraphs in each bag and hence can be done in $O(k^{2}t)$ time where $k=2^{4(t+1)}$. Therefore, a dual support of treewidth $2^{4(t+1)}$ can also be computed in time $\textrm{poly}(n^{t})$ since $|\mathcal{H}|\leq n^{3t+3}$ by Theorem 32.
Finally, the construction of an intersection support follows the construction of primal and dual supports. By the arguments above, it follows that a $\mathcal{K}$-easy tree decomposition can be computed in time $\textrm{poly}(n^{t})$ where the width of the resulting tree decomposition is $O(2^{t})$. Then a $2^{O(2^{t})}$-sparse tree decomposition can be computed in time $\textrm{poly}(n^{2^{t}})$. Hence, the total running time to compute an intersection support is $\textrm{poly}(n^{2^{t}})$.
@@ 7 Lower Bounds
Let $m=\lceil n/2\rceil$ and $N=2^{m}$. We construct a graph $G$ and non-piercing subgraphs $\mathcal{H}$ as follows. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of vertices colored $\mathbf{b}$ such that there is no edge between the vertices of $B$. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices colored $\mathbf{r}$. Let $\textrm{bin}(x)$ denote the binary representation of $x\in\mathbb{N}$. By $\textrm{bin}(x)$ of an indexed set $X$, we mean the subset of $X$ with indices corresponding to $1$s in $\textrm{bin}(x)$. Since $\log N=m$, for any integer $0\leq x\leq N$, $\textrm{bin}(x)$ of $C$ and $\textrm{bin}(x)$ of $R$ are well-defined.
The subsets $\mathcal{H}$ are indexed by pairs of vertices that are consecutive along a row, or along a column of the grid. For a subgraph $H_{i,j,i,j+1}$ that corresponds to a horizontal pair $e=\{b_{i,j},b_{i,j+1}\}$, we connect the end-points of $e$ to $\textrm{bin}(j)$ of $C$, $\textrm{bin}(j+1)$ of $C$ and $\textrm{bin}(i)$ of $R$. For a subgraph $H_{i^{\prime},j^{\prime},i^{\prime}+1,j^{\prime}}$ corresponding to a vertical pair $e^{\prime}=\{b_{i^{\prime},j^{\prime}},b_{i^{\prime}+1,j^{\prime}}\}$ we connect the end-points of $e^{\prime}$ to $\textrm{bin}(i^{\prime})$ of $R$, $\textrm{bin}(i^{\prime}+1)$ of $R$, and $\textrm{bin}(j^{\prime})$ of $C$.
@@
The subgraphs in $\mathcal{H}$ are connected. Let $H,H^{\prime}\in\mathcal{H}$. Since they correspond to distinct pairs of points of the grid, they differ in at least one grid vertex, and they also differ in their adjacency in either $R$ or in $C$. Hence, $H\setminus H^{\prime}$ is connected via the grid vertex in $H\setminus H^{\prime}$. Similarly, $H^{\prime}\setminus H$ is connected. Therefore, $\mathcal{H}$ is a collection of non-piercing subgraphs.
Each $H\in\mathcal{H}$ contains exactly two vertices of $\mathbf{b}$, which should be adjacent in any primal support for $(G,\mathcal{H})$. Therefore, any primal support $Q$ must contain a grid graph on $B$. Hence, $\textsc{tw}(Q)\geq N$. ∎
@@
The construction is similar to the construction of the primal system. Let $m=\lceil n/2\rceil$ and $N=2^{m}$. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of vertices. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices, none of which are pairwise adjacent.
The graph system $(G,\mathcal{H})$ is constructed as follows. A face $F_{i,j,i+1,j+1}$ for $i,j\in\{1,\ldots,N-1\}$ is the subset $\{b_{i,j},b_{i,j+1},b_{i+1,j},b_{i+1,j+1}\}$ of $B$. For each $F_{i,j,i+1,j+1}$ we add the subgraph $H_{i,j,i+1,j+1}$ consisting of the vertices of $F_{i,j,i+1,j+1}$, and we connect the vertices of the face to $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$ and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. Finally, for every two consecutive vertices on the grid, we put a vertex in the middle. Let $B^{\prime}$ be the set of vertices added. Thus, $B\cup B^{\prime}$ forms a grid of size $(2N-1)\times(2N-1)$. We add the vertices of $B^{\prime}$ to each subgraph containing the corresponding pair of vertices in its neighbour. For each subgraph $H$ containing a $b^{\prime}\in B^{\prime}$, we make $b^{\prime}$ adjacent to the vertices in $H\cap R$ and $H\cap C$. This results in connected induced subgraphs $\mathcal{H}$ of $G$.
@@
Since each vertex in $B^{\prime}$ lies in exactly two subgraphs, the subgraphs in $\mathcal{H}$ form a grid of size $(N/2)\times(N/2)$. Therefore, for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq N/2$. ∎
@@ 8 Conclusion
Based on the above discussion, one can think of the following question: Outerplanar graphs have treewidth at most $2$. If $G$ is a graph of treewidth $2$ and the subgraphs are non-piercing, is it possible to get a primal/dual or intersection support that has treewidth at most $2$?

### v2

@@ 6.4 Implementation
We show that $|\mathcal{H}|=O(n^{3t+3})$. The bound on the size of $\mathcal{K}$ follows analogously.
First, we will show that the VC-dimension of the set system defined by $(V(G),\mathcal{H})$ is at most $33t+$ 3. Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. We will show that no set of size more than $33t+$ 3 can be shattered. Suppose not. Let $S$ be a set with $43t+$ 4 vertices that can be shattered. There is an edge $\{x,y\}$ in $T$ such that the union of bags in $T_{x}$ contains a subset $S_{1}\subset S$ of at least $t+2$ vertices, and the union of bags in $T\setminus T_{x}$ contains a subset $S_{2}\subset S$ of at least $t+2$ vertices where $S_{1}\cap S_{2}=\emptyset$. Note that such subsets should exist since each bag contains at most $t+1$ vertices and $|S|=3t+4$.
There are at least $k=\binom{t+2}{\lfloor\frac{t+2}{2}\rfloor}$ subsets of $S_{i}$ of size $\lfloor\frac{t+2}{2}\rfloor$ for $i\in\{1,2\}$. Let $H_{1},H_{2},\ldots H_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{1}$, and $H^{\prime}_{1},H^{\prime}_{2},\ldots H^{\prime}_{k}$ be the distinct subgraphs in $\mathcal{H}$ defined by these $k$ sets in $S_{2}$. Note that $H_{i}$’s are pairwise incomparable in the containment order. Similarly, $H^{\prime}_{i}$ ’s are pairwise incomparable.
@@
Let $G$ be an $n$-vertex graph of treewidth $t$ and $\mathcal{H},\mathcal{K}$ be non-piercing subgraphs of $G$. Then, a primal and a dual support of treewidth $O(2^{t})$ and $O(2^{4t})$ respectively can be computed in time $\textrm{poly}(n^{t})$. And, an intersection support of treewidth at most $2^{O(2^{t})}$ can be computed in time $\textrm{poly}(n^{2^{t}})$.
@@
Since $\textsc{tw}(G)=t$, a tree decomposition $(T,\mathcal{B})$ of width $t$ can be computed in time $2^{O(t)}n$ where $T$ is a binary tree and has $O(n)$ number of nodes [bodlaender1998partial, bodlaender2016, BODLAENDER1996358].
For the construction of a primal support, we require an easy tree decomposition of $G$. We did a post-order traversal of the edges in $T$ and chose a minimal subgraph in $\mathcal{M}_{S}$ for each non-empty subset $S$ of an adhesion set $A$ in $T$. Since $A$ has at most $t+1$ vertices, there are at most $2^{t+1}$ choices for $S$. Also, by Theorem 34, $|\mathcal{H}|=n^{3t+3}$. Therefore, a primal support of treewidth $2^{t+2}+t$ can be computed in time $\textrm{poly}(n^{t})$.
For a dual support, we first construct a graph system that is $2^{4(t+1)}$-sparse. For such a construction, we again do a post-order traversal of the edges in $T$ and choose a subset $S$ of an adhesion set $A$ to select a pusher in $\mathcal{M}_{S}$. In $O(|\mathcal{H}|^{2})$ time, we can do the pushing operation at a subset $S$ to get new subgraphs. Therefore all the adhesion sets can be processed in time $O(2^{t}n|\mathcal{H}|^{2})$. Once a $2^{4(t+1)}$-sparse tree decomposition is obtained, the algorithm $k$-SDS puts a complete graph on the subgraphs in each bag and hence can be done in $O(k^{2}t)$ time where $k=2^{4(t+1)}$. Therefore, a dual support of treewidth $2^{4(t+1)}$ can also be computed in time $\textrm{poly}(n^{t})$ since $|\mathcal{H}|\leq n^{3t+3}$ by Theorem 34.
Finally, the construction of an intersection support follows the construction of primal and dual supports. By the arguments above, it follows that a $\mathcal{K}$-easy tree decomposition can be computed in time $\textrm{poly}(n^{t})$ where the width of the resulting tree decomposition is $O(2^{t})$. Then a $2^{O(2^{t})}$-sparse tree decomposition can be computed in time $\textrm{poly}(n^{2^{t}})$. Hence, the total running time to compute an intersection support is $\textrm{poly}(n^{2^{t}})$. ∎
@@ 7 Lower Bounds
Let $m=\lceil n/2\rceil$ and $N=2^{m}$. We construct a graph $G$ and non-piercing subgraphs $\mathcal{H}$ as follows. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices colored $\mathbf{b}$ such that there is no edge between the vertices of $B$. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices colored $\mathbf{r}$. Let $\textrm{bin}(x)$ denote the binary representation of $x\in\mathbb{N}$. By $\textrm{bin}(x)$ of an indexed set $X$, we mean the subset of $X$ with indices corresponding to $1$s in $\textrm{bin}(x)$. Since $\log N=m$, for any integer $0\leq x\leq N$, $\textrm{bin}(x)$ of $C$ and $\textrm{bin}(x)$ of $R$ are well-defined.
For a horizontal pair of grid vertices $e=\{b_{i,j},b_{i,j+1}\}$, we connect these vertices to $\textrm{bin}(j)$ of $C$, $\textrm{bin}(j+1)$ of $C$ and $\textrm{bin}(i)$ of $R$. This gives an induced subgraph $H_{i,j,i,j+1}$. Similarly, for a vertical pair $e^{\prime}=\{b_{i^{\prime},j^{\prime}},b_{i^{\prime}+1,j^{\prime}}\}$, we connect these vertices to $\textrm{bin}(i^{\prime})$ of $R$, $\textrm{bin}(i^{\prime}+1)$ of $R$, and $\textrm{bin}(j^{\prime})$ of $C$. This gives us an induced subgraph $H_{i^{\prime},j^{\prime},i^{\prime}+1,j^{\prime}}$. This completes simultaneously, the construction of $G$ and the subgraphs $\mathcal{H}$.
@@
The subgraphs in $\mathcal{H}$ are connected. Let $H,H^{\prime}\in\mathcal{H}$ be arbitrary. Since they correspond to distinct pairs of points of the grid, they differ in at least one grid vertex, and they also differ in their adjacency in either $R$ or in $C$. Hence, $H\setminus H^{\prime}$ is connected via the grid vertex in $H\setminus H^{\prime}$. Similarly, $H^{\prime}\setminus H$ is connected. Therefore, $\mathcal{H}$ is a collection of non-piercing subgraphs of $G$.
Each $H\in\mathcal{H}$ contains exactly two vertices of $B$, which should be adjacent in any primal support for $(G,\mathcal{H})$. Therefore, any primal support $Q$ must contain a grid graph on $B$. Hence, $\textsc{tw}(Q)\geq N$.
@@
The construction is similar to the construction of the primal system. Let $m=\lceil n/2\rceil$ and $N=2^{m}$. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices, none of which are pairwise adjacent.
The graph system $(G,\mathcal{H})$ is constructed as follows. A face $F_{i,j,i+1,j+1}$ for $i,j\in\{1,\ldots,N-1\}$ is the subset $\{b_{i,j},b_{i,j+1},b_{i+1,j},b_{i+1,j+1}\}$ of $B$. For each $F_{i,j,i+1,j+1}$, we connect all its vertices to $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. Finally, for every two consecutive vertices on the grid, we put a vertex in the middle. Let $B^{\prime}$ be the set of vertices added. Thus, $B\cup B^{\prime}$ forms a grid of size $(2N-1)\times(2N-1)$. For each vertex of $B^{\prime}$ that lies between two consecutive vertices $x,y$ of $B$ row-wise or column-wise, we make it adjacent to all the vertices in $R\cup C$ that are adjacent to both $x$ and $y$. This completes the construction of $G$. Now, for each face $F_{i,j,i+1,j+1}$, we add an induced subgraph $H_{i,j,i+1,j+1}$ to $\mathcal{H}$ consisting of the vertices of $F_{i,j,i+1,j+1}$, the four vertices of $B^{\prime}$ that lie between the vertices of this face, and the common neighbours of these 8 vertices in $R\cup C$ i.e., $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. This results in connected induced subgraphs $\mathcal{H}$ of $G$.
@@
Since each vertex in $B^{\prime}$ lies in exactly two subgraphs, the subgraphs in $\mathcal{H}$ form a grid of size $(N-1)\times(N-1)$. Therefore, for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq N-1$.
@@ 8 Conclusion
From our results, it follows that a non-piercing outerplanar graph system admits an outerplanar support. Since $\textsc{tw}(G)=2$ for an outerplanar graph $G$, we wonder if for a non-piercing graph system of treewidth 2, there exists a primal/dual or intersection support of treewidth 2.